OpenStudy (anonymous):

One circular table has a diameter of 9 ft, and another circular table has a diameter of 14 ft. How much greater is the area of the larger table? Round to the nearest whole number.

OpenStudy (jmark):

Area Formula for a Circle is: A = pi * r^2 pi * (4.5)^2 = pi*20.25 pi * 7^2 = pi*49 = 49pi-20.25pi = 90.32 The Larger table is 90.32 sq ft bigger than the smaller table. Maths Questions for Class 7 ICSE - http://scoremore.beep.com/maths-questions-for-class-7-icse-2014-07-25-1.htm